OpenStudy (anonymous):

Solve the system. What is the value of x? y = 3x and 4x + 2y = 30 A. 3 B. -3 C. 9 D. -9

OpenStudy (anonymous):

Substitute y = 3x in second equation.

OpenStudy (anonymous):

3

OpenStudy (ash2326):

y=3x 4x+2y=30 put y=3x in second equation 4x+2*3x=30 10x=30 x=3

OpenStudy (mathteacher1729):

(equation ONE) y = 3x and (equation TWO) 4x + 2y = 30 Substitute y = 3x into the 2nd equation: 4x + 2(3x) = 30 simplify 4x + 6x = 30 10x = 30 x = 3 Now that you know x = 3, you can put that into the second equation 4x + 2y = 30 4(3) + 2y = 30 12 + 2y = 30 subtract 12 from both sides 2y = 18 divide by 2 y = 9 =========== So x = 3, y = 9. =========== You can see this by going to http://www.wolframalpha.com and typing plot y = 3x and 4x + 2y = 30 (note the intersection of the two lines is at (3,9). :)